OpenStudy (anonymous):

3y = –1/2x + 2 y = –x + 9

OpenStudy (anonymous):

Solve by substitution. !~

OpenStudy (anonymous):

y=-x+9 thus 3*(-x+9)=-1/2x+2 -3x+27=-1/2x+2 5/2x=25 x=25*2/5 x=10

jhonyy9 (jhonyy9):

in the first equation x is in denominator or numerator ?